OpenStudy (anonymous):

I need help with factorizing quadratic trinomials. I have a math test tomorrow, and the math book's explanation doesn't make any sense at all. Could someone explain to me how to go about factorizing quadratic trinomials? Some problems: x^2+9x+20 x^2-2x-8 3x^2-6x-24 -4+4x-x^2

OpenStudy (anonymous):

you need to tackle them one by one the easy one first

Parth (parthkohli):

First two and last two terms. Just for a little sample, I'd do the first. You have \(\rm 2a + 2 +a b + b\) First two - \(\rm 2(a + 1)\) Last two - \(\rm b(a + 1)\) You're left with \(\rm 2(a + 1) + b(a + 1) = (2 + b)(a + 1)\)

OpenStudy (anonymous):

@ParthKohli Right, thanks. What about, say, \[x^2+9x+20\]?

OpenStudy (anonymous):

4x5=20 use that

OpenStudy (anonymous):

also 4+5 =9

OpenStudy (anonymous):

x^2+9x+20=0 (x+4)(x+5)=0 check them by mult n add x^2 +4x+5x+20=0 x^2+9+20=0 checked

OpenStudy (anonymous):

x^2-2x-8=0 hint -4(2)=-8 and -4+2=-2

OpenStudy (anonymous):

can you try that?

OpenStudy (anonymous):

Alright, got it. Thanks! :)

OpenStudy (anonymous):

ok YW good luck :D